CLASSICAL STUDY OF ENERGY FLOW IN HEI_{2}

\underline{R. C. MAYRHOFER}, R. J. KOVACICH, Department of Physical Science, Kutztown University, Kutztown, PA 19530.

The classical dynamics of energy flow within T-shaped HeI_{2} is studied by numerically integrating the appropriate equations of motion. From this two dimensional study, a relationship is developed between the exchange of energy in the van der Waal degree of freedom and the initial phase of I_{2}. The breakdown in phase space is also examined by constructing Poincarè surface of section plots for various initial conditions. We show that these plots can also be generated from a perturbative formalism using the Morse oscillator as our zero order picture.